The length of the sides of a trapezoid are (x+2), (2x+5) and two sides of (2x-1). Find the length of the shortest side if the perimter is 82.
Answer by robertb(5830)

x+2 +2x + 5 + 2(2x-1) = 82.
3x+7 + 4x-2 = 82,
7x + 5 = 82,
7x = 77,
x = 11.
x+2 = 13,
2x + 5 = 27,
2x-1 = 21.
so now you know which is the shortest side!